### v2

@@ 1 Introduction
  The gold standard for treatment effect estimation remains the RCT. The random treatment allocation in the study guards against the effect of confounding As a result, under standard identifiability assumptions, we can obtain an unbiased estimate of the treatment effect. There are, however, drawbacks to the use of RCTs (Frieden, 2017), such as the high financial cost or the limited sample sizes. While they may be powered enough to detect the Average Treatment Effect (ATE), they are typically underpowered for subgroup effects. Furthermore, strict eligibility criteria and selection bias may lead to a sample non-representative of the target population (National Academies of Sciences, Engineering, and Medicine, 2022).
Observational studies provide an alternative source of data for treatment effect estimation. In contrast to RCTs, they generally benefit from larger sample sizes and tend to be more representative of the target population. However, observational studies may suffer from unobserved confounding, i.e. unobserved variables that affect both the outcome of interest and the treatment allocation, and can lead to biased treatment effect estimates. For example, a patient’s underlying health status might influence both the treatment they receive and their eventual outcome, yet remain unmeasured.
  The complementary nature of RCT and observational data indicates the potential benefits from combining the two sources to obtain better heterogeneous treatment effect estimates. Such data fusion approaches have received increased interest in recent years. A range of methods have been developed to integrate information from multiple sources of data, including a limited number targeting causal inference problems (Colnet et al. (2023), Lin et al. (2025b)). These typically rely on strong, untestable assumptions about the structure of the confounding effect and, moreover, are often not able to provide uncertainty quantification around the point estimates of heterogeneous treatment effects.
@@ 2 Related Literature
Heterogeneous treatment effect (HTE) estimation has received renewed interest in recent years. The most common strategy is to focus on the conditional average treatment effect (CATE) function, which quantifies the expected treatment effect given particular covariate values. Under suitable identifiability assumptions (Dahabreh and Hernán, 2019), the CATE can be estimated parametrically - e.g. using linear regression - or non-parametrically, e.g. using nearest-neighbour matching, kernel methods (Brantner et al., 2023), or tree-based methods including causal random forests (Wager and Athey, 2018) and their Bayesian counterpart (Hahn et al., 2020).
@@
Data fusion methods for CATE estimation, the focus of this paper, follow three broad lines of work. One strategy is to use experimental data to debias observational estimates by learning a correction or bias function. Kallus et al. (2018) introduce a two-step approach that corrects hidden confounding in the absence of covariate overlap, relying on the strong assumption of linear confounding and the ability to identify this correction term parametrically. Yang et al. (2022b) relax this linearity assumption and provide identifiability and efficiency results under more flexible structural models. Hatt et al. (2022) develop a representation-learning strategy that learns shared features and confounded outcome models from observational data, using the RCT to estimate a bias function that debiases these models, with finite-sample bounds that highlight how performance depends on sample size, distribution shift, and bias complexity. Wu and Yang (2022) propose an R-learner that achieves consistency and asymptotic efficiency under covariate overlap between the RCT and the observational study. A second direction consists of “test-then-pool” strategies, in which the trial and observational datasets are combined only if the observational study appears sufficiently unbiased. For instance, Yang et al. (2023) implement such a strategy, although their approach is limited to linear treatment effect models and relies on strong parametric assumptions. A third line of work uses Bayesian dynamic borrowing to regulate how much information is extracted from the observational study. For instance, Lin and Evans (2023) propose a Bayesian dynamic borrowing approach through a power likelihood to reduce the effect of confounding and control the degree of information we borrow from the observational study.
Our method, Causal-ICM, takes a Bayesian dynamic borrowing approach based on multi-task GPs, whereby each potential outcome is treated as a distinct task. Multi-task GPs have been successfully employed for causal inference using solely observational data (Alaa and van der Schaar, 2017) and extended to handle hierarchical hidden confounders (Witty et al., 2020). Others have built on this framework to develop a general class of counterfactual multi-task deep kernel models that efficiently estimate causal effects and learn policies by stacking coregionalized GPs and deep kernels (Caron et al., 2022). GPs have also been employed as a matching tool for causal inference (Huang et al., 2023). However, these methods are limited to purely observational settings; here we extend these ideas by enabling inference in a data-fusion setting. Our approach is most similar to concurrent work employing multi-task GPs for pseudo-outcome regression to obtain error bounds on the observational bias term (Fawkes et al., 2025).
@@ 3 Methodology
3.1 Data Fusion for Heterogenous Treatment Effect Esimation
In this work, we focus on combining information from an observational dataset and a RCT (experimental) dataset. We begin with some notation. Let $S\in\{\text{o},\text{e}\}$ indicate the study, with ‘e’ denoting the experimental sample and ‘o’ the observational one. Let $A\in\{0,1\}$ be the binary treatment indicator. We observe the same baseline covariates in both studies, so each individual has $\mathbf{X}\in\mathcal{X}^{S}\subseteq\mathbb{R}^{p}$, where $\mathcal{X}^{\text{e}}$ and $\mathcal{X}^{\text{o}}$ overlap but $\mathcal{X}^{\text{o}}$ is not necessarily contained in $\mathcal{X}^{\text{e}}$. Without loss of generality we divide the baseline covariates into three subsets, the treatment effect modifiers $\mathbf{X}_{\tau}\in\mathcal{X}_{\tau}^{\text{e}}$, the confounders, $\mathbf{X}_{W}\in\mathcal{X}_{W}^{\text{e}}$, and the selection variables, $\mathbf{X}_{S}\in\mathcal{X}_{S}^{\text{e}}$, such that $\mathbf{X}=\mathbf{X}_{\tau}\cup\mathbf{X}_{W}\cup\mathbf{X}_{S}$ Finally, we observe a continuous outcome $Y\in\mathbb{R}$. The counterfactual outcome $Y_{i}^{a}$ denotes the value individual $i$ would have had under treatment $A=a$. The causal estimand of interest is the CATE comparing $A=1$ versus $A=0$ in the target population given the baseline covariates X{X}, defined as
@@ 3.1 Data Fusion for Heterogenous Treatment Effect Esimation
for $\mathbf{x}_{\tau}\in\mathcal{X}^{\text{e}}\cup\mathcal{X}^{\text{o}}$. The difference in the conditional expectations of the outcomes in each study is
@@
    Mean conditional exchangeability over treatment for the treatment contrast given $\mathbf{X}_{\tau}$ in the RCT
@@
    Mean conditional exchangeability over selection into the RCT for the treatment contrast given $\mathbf{X}_{\tau}$, $\mathbf{X}_{S}$:
    $\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S}]=\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\text{e}]$
@@
Under assumptions A1 - A5 the CATE can be identified from the RCT data as
$\tau(\mathbf{x}_{\tau})=\omega^{\text{e}}(\mathbf{x}_{\tau})$
for $\mathbf{x}_{\tau}\in\mathcal{X}_{\tau}^{\text{e}}$. The derivation is provided in the Appendix A. For notational simplicity, and with a slight abuse of notation, we henceforth use $\mathbf{X}$ and $\mathbf{x}$ to denote any subset of covariates, whether treatment-effect modifiers or selection covariates.
@@ 3.2 Multi-task Gaussian Processes
For $\mathbf{f}(\mathbf{x})=(f^{\text{e}}\left(\mathbf{x}\right),f^{\text{o}}\left(\mathbf{x}\right))$, we assign a multi-task GP prior $\textbf{f}\sim\mathcal{GP}\left(\boldsymbol{\mu},K\right)$, where $\boldsymbol{\mu}$ is a bivariate mean function (taken again to be 0) and $K$ is a positive $2\times 2$ matrix-valued covariance function that maps input points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ to matrices quantifying the relationship of the outputs at these inputs. We specify $K$ in the next section. The key strength of multi-task GP s is the ability to share information between tasks, which we will leverage for data fusion. See Appendix B Fig 3 for an illustrative comparison to independent GP s.
@@ 3.3 Intrinsic coregionalisation models (ICMs)
where $u_{q}\left(\mathbf{x}\right)$ are independent zero-mean latent functions taken from the same scalar GP, $u_{q}\sim\mathcal{GP}\left(0,k\right)$ for a chosen kernel $k$. The scalar coefficients $(\alpha_{q}^{\text{e}},\alpha_{q}^{\text{o}})$ are which will be used to construct the coregionalization matrix defined shortly, which governs the dependence structure. The rank of the ICM $Q$ is the number of independent components, which we set to $Q=2$ for the single arm case; we provide a discussion on the interpretation of this later.
@@
The posterior distribution over $f^{\text{e}}$ can then be computed using standard theory (see Appendix C). For a given test point $\mathbf{x}_{*}$, we have that the posterior distribution is Gaussian
@@ 3.4 Causal-ICM
We now introduce Causal-ICM. Specifically, we propose an interpretable parametrisation of the ICM for causal inference and introduce a bespoke procedure for learning its coefficients, rather than relying on standard marginal likelihood maximisation. This serves to control the influence of the observational dataset and avoid “swamping” the RCT data. Causal-ICM consists of the following choices for the ICM coefficients:
@@
Under this parameterization, $\rho\in(0,1)$ controls the degree of borrowing, with $\rho\to 1$ corresponding to maximum borrowing and $\rho\to 0$ to none (We provide an additional interpretation of $\rho$ in Appendix D, Proposition D.1). Although this implies $\beta^{\text{e}}=\beta^{\text{o}}=1$ (and $\beta_{\text{eo}}=\rho$), the model remains flexible as long as the kernel $k(\mathbf{x},\mathbf{x})$ includes an independent scaling term (e.g., the variance in an RBF kernel) and the scales of $f^{\text{e}}$ and $f^{\text{o}}$ are similar.
@@
Thus, $f^{\text{e}}=u_{1}$ and $f^{\text{o}}$ is a scaled version of $f^{\text{e}}$ plus a term capturing confounding. Choosing a positive $\rho\in(0,1)$ is reasonable, as $f^{\text{o}}(\mathbf{x})$ is expected to be highly correlated with $f^{\text{e}}(\mathbf{x})$ - both model the conditional expectation of the outcome - with $f^{\text{o}}$ additionally contaminated by confounding. In practice, restricting $\alpha_{2}^{\text{e}}$ minimally affects flexibility while improving interpretability. Remaining hyperparameters, including kernel parameters and sampling variance, are estimated via marginal likelihood maximisation.
@@ 3.4.1 Variance bound for conditional mean function under Causal-ICM
When evaluating treatment effects, it is important to quantify uncertainty alongside point estimates. The multi-task Gaussian process naturally provides this uncertainty via the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$. In data fusion, a key concern is that large observational samples may “swamp” the experimental data, yielding biased and overconfident estimates of $\tau(\mathbf{x})$. Causal-ICM mitigates this, as shown by a lower bound on the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$ even as $n^{\text{o}}$ grows. This variance bound can also be seen as formally limiting the information flow from the observational dataset, similar to cut models (Lin et al., 2025a).
@@
Let $n^{\textnormal{e}},n^{\textnormal{o}}\geq 0$ denote the sizes of the experimental and observational datasets respectively. We outline the proof for $n^{\textrm{e}}=0$ here, which corresponds to the prior-only case for the RCT. The proof for the general case $n^{\textrm{e}}>0$ is deferred to Appendix E. The key is that when $n^{\textrm{e}}=0$, the posterior variance can be written as
@@
An intuitive understanding of the result is as follows: for two random variables with correlation $\rho$, knowledge of one random variable does not inform us of the exact value of the other unless $\rho=1$. The key interpretation of Proposition 3.1 in the data fusion context is as follows: even as $n^{\text{o}}\to\infty$, the posterior variance of $f^{\text{e}}(\mathbf{x}_{*})$ conditional on the full dataset will at most decrease by a factor of $(1-\rho^{2})$ relative to the posterior variance given the experimental data only, where $|\rho|\leq 1$. We conjecture that this inequality is tight as $n^{\text{o}}\to\infty$. This can be interpreted as limiting the ‘effective sample size’ of the observational dataset to be $1/\left(1-\rho^{2}\right)$ relative to the experimental dataset. Crucially, this variance bound protects from having very tight credible intervals centred around a biased estimate - the posterior uncertainty will accurately reflect our distrust of the observational dataset. We also expect this ‘effective sample size’ effect to hold for the posterior mean of $f^{\text{e}}(\mathbf{x}_{*})$, where there the bias introduced to the posterior mean by $\mathcal{D}_{\text{o}}$ is limited again by the choice of $\rho$. We leave this interesting direction for future work. Although our focus is on the conditional mean functions, our approach also provides the posterior distribution of the confounding function, offering additional insights that may be of independent interest (see Appendix F).

### v3

@@ 1 Introduction
  The gold standard for treatment effect estimation remains the RCT. The random treatment allocation in the study guards against the effect of confounding As a result, under standard identifiability assumptions, we can obtain an unbiased estimate of the treatment effect. However, RCTs have important drawbacks (Frieden, 2017), including high financial costs or the limited sample sizes. While often powered to detect the Average Treatment Effect (ATE), they are typically underpowered for subgroup effects. Furthermore, strict eligibility criteria and selection bias can limit representativeness of the target population (National Academies of Sciences, Engineering, and Medicine, 2022).
  Observational studies offer an alternative source of data for treatment effect estimation, typically benefiting from larger sample sizes and better representativeness of the target population. However, they may be susceptible to unobserved confounding - unmeasured variables that affect both outcomes and treatment assignment - leading to biased treatment effect estimates. For example, a patient’s underlying health status might influence both the treatment they receive and their eventual outcome, yet remains unmeasured.
  The complementary nature of RCT and observational data indicates the potential benefits from combining the two sources to obtain better heterogeneous treatment effect estimates. Such data fusion approaches have received increased interest in recent years. A range of methods have been developed to integrate information from multiple sources of data, including a limited number targeting causal inference problems (Colnet et al. (2024), Lin et al. (2024)). These typically rely on strong, untestable assumptions about the structure of the confounding effect and, moreover, are often not able to provide uncertainty quantification around the point estimates of heterogeneous treatment effects.
@@ 2 Related Literature
Heterogeneous treatment effect (HTE) estimation has received renewed interest in recent years. The most common strategy is to focus on the conditional average treatment effect (CATE) function, which quantifies the expected treatment effect given particular covariate values. Under suitable identifiability assumptions (Dahabreh and Hernán, 2019), the CATE can be estimated parametrically - e.g. using linear regression - or non-parametrically, e.g. using nearest-neighbour matching, kernel methods (Brantner et al., 2023), or tree-based methods including causal random forests (Wager and Athey, 2018) and their Bayesian counterpart (Hahn et al., 2020). A related strand of work studies semi-parametric and debiased machine learning approaches that combine flexible outcome models with orthogonal score constructions to obtain robust estimates under high-dimensional confounding.
@@
Data fusion methods for CATE estimation, the focus of this paper, follow three broad lines of work. One strategy is to use experimental data to debias observational estimates by learning a correction or bias function. Kallus et al. (2018) introduce a two-step approach that corrects hidden confounding in the absence of covariate overlap, relying on the strong assumption of linear confounding and the ability to identify this correction term parametrically. Yang et al. (2025b) relax this linearity assumption and provide identifiability and efficiency results under more flexible structural models. Hatt et al. (2022) develop a representation-learning strategy that learns shared features and confounded outcome models from observational data, using the RCT to estimate a bias function that debiases these models, with finite-sample bounds that highlight how performance depends on sample size, distribution shift, and bias complexity. Wu and Yang (2022) propose an R-learner that achieves consistency and asymptotic efficiency under covariate overlap between the RCT and the observational study. A second direction consists of “test-then-pool” strategies, in which the trial and observational datasets are combined only if the observational study appears sufficiently unbiased. For instance, Yang et al. (2023) implement such a strategy, although their approach is limited to linear treatment effect models and relies on strong parametric assumptions. A third line of work uses Bayesian dynamic borrowing to regulate how much information is extracted from the observational study. For instance, Lin et al. (2025) propose a Bayesian dynamic borrowing approach through a power likelihood to reduce the effect of confounding and control the degree of information we borrow from the observational study.
Our method, Causal-ICM, takes a Bayesian dynamic borrowing approach based on multi-task GPs, whereby each potential outcome is treated as a distinct task. Multi-task GPs have been successfully employed for causal inference using solely observational data (Alaa and van der Schaar, 2017) and extended to handle hierarchical hidden confounders (Witty et al., 2020). Others have built on this framework to develop a general class of counterfactual multi-task deep kernel models that efficiently estimate causal effects and learn policies by stacking coregionalized GPs and deep kernels (Caron et al., 2022). GPs have also been employed as a matching tool for causal inference (Huang et al., 2023). Recent work has also explored Gaussian-process-based partially linear models for heterogeneous treatment effects, combining parametric effect components with flexible GP nuisance estimation (Horii and Chikahara, 2024). However, these approaches focus on single-source observational settings. In contrast, our work develops a multi-task GP framework specifically designed for causal data fusion between experimental and observational sources. Our approach is most similar to concurrent work employing multi-task GPs for pseudo-outcome regression to obtain error bounds on the observational bias term (Fawkes et al., 2025).
@@ 3 Methodology
3.1 Data Fusion for Heterogeneous Treatment Effect Estimation
In this work, we focus on combining information from an observational dataset and a RCT (experimental) dataset. We begin with some notation. Let $S\in\{\text{o},\text{e}\}$ indicate the study, with ‘e’ denoting the experimental sample and ‘o’ the observational one. Let $A\in\{0,1\}$ be the binary treatment indicator. We observe the same baseline covariates in both studies, so each individual has $\mathbf{X}\in\mathcal{X}^{S}\subseteq\mathbb{R}^{p}$, where $\mathcal{X}^{\text{e}}$ and $\mathcal{X}^{\text{o}}$ overlap but $\mathcal{X}^{\text{o}}$ is not necessarily contained in $\mathcal{X}^{\text{e}}$. Our methodology requires overlap between the covariate distributions of the RCT and the observational study, but it does not require the support of the RCT covariates to be fully contained within that of the observational study. In regions with little or no randomized evidence, the observational study may still provide useful information for generalising treatment-effect estimates, although uncertainty should increase accordingly.
Without loss of generality, we distinguish three possible roles for baseline covariates: treatment-effect modifiers $\mathbf{X}_{\tau}\in\mathcal{X}_{\tau}^{\text{e}}$, confounders $\mathbf{X}_{W}\in\mathcal{X}_{W}^{\text{e}}$, and selection variables $\mathbf{X}_{S}\in\mathcal{X}_{S}^{\text{e}}$, such that $\mathbf{X}=\mathbf{X}_{\tau}\cup\mathbf{X}_{W}\cup\mathbf{X}_{S}$. Here, confounders are variables that act as common causes of treatment and outcome, treatment-effect modifiers are variables indexing the heterogeneity of interest, and selection variables are variables whose distribution may differ between the RCT and the target population. These roles are not assumed to be mutually exclusive. Finally, we observe a continuous outcome $Y\in\mathbb{R}$. The counterfactual outcome $Y_{i}^{a}$ denotes the value individual $i$ would have had under treatment $A=a$. The causal estimand of interest is the CATE in the target population given the baseline covariates X{X}, defined for $\mathbf{x}_{\tau}\in\mathcal{X}^{\text{e}}\cup\mathcal{X}^{\text{o}}$ as
@@ 3.1 Data Fusion for Heterogeneous Treatment Effect Estimation
The difference in the conditional expectations of the outcomes in each study is denoted
@@
    Mean conditional exchangeability over treatment for the treatment contrast given $\mathbf{X}_{\tau}$ in the RCT:
@@
    Mean conditional exchangeability over selection into the RCT for the treatment contrast given $\mathbf{X}_{\tau}$, $\mathbf{X}_{S}$: $\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S}]=\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\text{e}]{\color[rgb]{0,1,0}}$
@@
  6. A6
    Conditional invariance of selection variable distributions between the RCT and the target population: $P(X_{S}|X_{\tau})=P(X_{S}|X_{\tau},S=\text{e})$.
Under assumptions A1 - A6 the CATE can be identified from the RCT data as $\tau(\mathbf{x}_{\tau})=\omega^{\text{e}}(\mathbf{x}_{\tau})$ for $\mathbf{x}_{\tau}\in\mathcal{X}_{\tau}^{\text{e}}$. The derivation is provided in Appendix A, and a discussion on Assumptions A4 and A6 is provided on Appendix B. For notational simplicity, and with a slight abuse of notation, we henceforth use $\mathbf{X}$ and $\mathbf{x}$ to denote whichever subset of covariates is relevant in context, including treatment-effect modifiers, confounders, or selection covariates.
@@ 3.2 Multi-task Gaussian Processes
For $\mathbf{f}(\mathbf{x})=(f^{\text{e}}\left(\mathbf{x}\right),f^{\text{o}}\left(\mathbf{x}\right))$, we assign a multi-task GP prior $\textbf{f}\sim\mathcal{GP}\left(\boldsymbol{\mu},K\right)$, where $\boldsymbol{\mu}$ is a bivariate mean function (taken again to be 0) and $K$ is a positive $2\times 2$ matrix-valued covariance function that maps input points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ to matrices quantifying the relationship of the outputs at these inputs. We specify $K$ in the next section. The key strength of multi-task GP s is the ability to share information between tasks, which we will leverage for data fusion. See Appendix C Fig 3 for an illustrative comparison to independent GP s.
@@ 3.3 Intrinsic coregionalisation models (ICMs)
where $u_{q}\left(\mathbf{x}\right)$ are independent zero-mean latent functions taken from the same scalar GP, $u_{q}\sim\mathcal{GP}\left(0,k\right)$ for a chosen kernel $k$. The scalar coefficients $(\alpha_{q}^{\text{e}},\alpha_{q}^{\text{o}})$ are which will be used to construct the coregionalization matrix defined shortly, which governs the dependence structure. The rank, $Q$, of the ICM is the number of independent components, which we set to $Q=2$ for the single arm case; we provide a discussion on the interpretation of this later.
@@
The posterior distribution over $f^{\text{e}}$ can then be computed using standard theory (see Appendix D). For a given test point $\mathbf{x}_{*}$, we have that the posterior distribution is Gaussian:
@@ 3.4 Causal-ICM
We now introduce Causal-ICM. Specifically, we propose an interpretable parametrisation of the ICM for causal inference and introduce a bespoke procedure for learning its coefficients, rather than relying on standard marginal likelihood maximisation. This serves to control the influence of the observational dataset and avoid a situation in which a large confounded observational sample yields biased estimates with artificially narrow uncertainty intervals relative to the RCT. Causal-ICM consists of the following choices for the ICM coefficients:
@@
Under this parameterization, $\rho\in(0,1)$ controls the degree of borrowing, with $\rho\to 1$ corresponding to maximum borrowing and $\rho\to 0$ to none (We provide an additional interpretation of $\rho$ in Appendix E, Proposition E.1). Although this implies $\beta^{\text{e}}=\beta^{\text{o}}=1$ (and $\beta_{\text{eo}}=\rho$), the model remains flexible as long as the kernel $k(\mathbf{x},\mathbf{x})$ includes an independent scaling term (e.g., the variance in an RBF kernel) and the scales of $f^{\text{e}}$ and $f^{\text{o}}$ are similar.
@@
Thus, $f^{\text{e}}=u_{1}$ and $f^{\text{o}}$ is a scaled version of $f^{\text{e}}$ plus a term capturing confounding. Setting $\alpha_{2}^{\text{e}}=0$ means that the experimental regression surface is modelled directly as a single GP, while the observational regression surface decomposes into a shared component and an observational-specific component. Choosing $\rho\in(0,1)$ is reasonable, as $f^{\text{o}}(\mathbf{x})$ is expected to be positively correlated with $f^{\text{e}}(\mathbf{x})$, both model the conditional expectation of the outcome, with $f^{\text{o}}$ additionally contaminated by confounding. In practice, restricting $\alpha_{2}^{\text{e}}$ minimally affects flexibility while improving interpretability. Remaining hyperparameters, including kernel parameters and sampling variance, are estimated via marginal likelihood maximisation.
@@ 3.4.1 Variance bound for conditional mean function under Causal-ICM
When evaluating treatment effects, it is important to quantify uncertainty alongside point estimates. The multi-task Gaussian process provides this directly via the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$. In data fusion, a key concern is that large observational samples may lead to biased estimates with unrealistically narrow uncertainty intervals when unmeasured confounding is present. Causal-ICM mitigates this, as shown by a lower bound on the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$ even as $n^{\text{o}}$ grows. This variance result can be viewed as formally limiting the information flow from the observational dataset, similar to cut models (Lin et al., 2025).
@@
Let $n^{\textnormal{e}},n^{\textnormal{o}}\geq 0$ denote the sizes of the experimental and observational datasets respectively. We outline the proof for $n^{\textrm{e}}=0$ here, which corresponds to the prior-only case for the RCT. The proof for the general case $n^{\textrm{e}}>0$ is deferred to Appendix F. The key is that when $n^{\textrm{e}}=0$, the posterior variance can be written as
@@
Proposition 3.1 shows that the Causal-ICM posterior variance is lower bounded by a non-degenerate fraction of the experimental-only posterior variance. The same decomposition also yields an upper bound. Writing
$V_{D}^{\mathrm{e}}(x^{*})=V_{D_{\mathrm{e}}}^{\mathrm{e}}(x^{*})-\rho^{2}\bigl(k(x^{*},X^{\mathrm{o}})-k^{\prime}(x^{*},X^{\mathrm{o}})\bigr)\Sigma_{eo}^{-1}\bigl(k(X^{\mathrm{o}},x^{*})-k^{\prime}(X^{\mathrm{o}},x^{*})\bigr),$
and noting that the quadratic form is non-negative, we obtain
$(1-\rho^{2})V_{D_{\mathrm{e}}}^{\mathrm{e}}(x^{*})\leq V_{D}^{\mathrm{e}}(x^{*})\leq V_{D_{\mathrm{e}}}^{\mathrm{e}}(x^{*}).$
Hence, incorporating observational data cannot inflate the posterior variance beyond the RCT-only baseline, nor arbitrarily concentrate the posterior when $\rho\in(0,1)$.
An intuitive understanding of the result is as follows: for two random variables with correlation $\rho$, knowledge of one random variable does not inform us of the exact value of the other unless $\rho=1$. The key interpretation of Proposition 3.1 in the data fusion context is as follows: even as $n^{\text{o}}\to\infty$, the posterior variance of $f^{\text{e}}(\mathbf{x}_{*})$ conditional on the full dataset will at most decrease by a factor of $(1-\rho^{2})$ relative to the posterior variance given the experimental data only, where $|\rho|\leq 1$. We conjecture that this inequality is tight as $n^{\text{o}}\to\infty$. This can be interpreted as limiting the ‘effective sample size’ of the observational dataset to be $1/\left(1-\rho^{2}\right)$ relative to the experimental dataset. Crucially, this variance bound protects from having very tight credible intervals centred around a biased estimate - the posterior uncertainty will accurately reflect our distrust of the observational dataset. We also expect this ‘effective sample size’ effect to hold for the posterior mean of $f^{\text{e}}(\mathbf{x}_{*})$, where there the bias introduced to the posterior mean by $\mathcal{D}_{\text{o}}$ is limited again by the choice of $\rho$. We leave this interesting direction for future work. Although our focus is on the conditional mean functions, our approach also provides the posterior distribution of the confounding function, offering additional insights that may be of independent interest (see Appendix G).